# Governing Together: Toward Infrastructure for Community-Run Social Media

### v1

@@ 6.1. Modularity, Polycentricity, and Forkability: Principles for Inter-community Governance
The ideas in our findings come together as a collection of resources, norms, and infrastructures that enable communities to collaborate and cooperate on governance issues from the bottom-up. In Figure 3, we sketch them as one ecosystem of inter-community governance on community-run social media. We present this illustration as a conceptual artifact that can prompt future design work aiming to support, shape, and scaffold dynamics of inter-community governance. Some of the concepts presented above suggest ideas that could be easily developed in the near future. However, as our findings note, the success of any system suggested by these ideas are likely to require shifts in broader social and cultural expectations around how communities operate on the Fediverse. For example, many workshop participants envisioned decentralized social media as consisting of “many fediverses,” an idea that they suggested was considered controversial by those who wanted to garner the benefits of scale. Rather than focusing on technical tools as solutions, we emphasize devising infrastructures that can create opportunities for communities to organize and reshape norms together. This builds on prior work from HCI that examines and intervenes in technical infrastructures — whether through creating new tools standardizing processes (e.g., decision-making) that produce new possibilities (or constraints) for users and communities broadly (Edwards et al., 2010) — as well as human infrastructures that do work to coordinate and manage these systems (Lee et al., 2006). Designing infrastructures also presents opportunities for discussing social norms and discussing what types of decisions should be made at a global versus local level (Edwards et al., 2009). As we note in §5.6, this may start as simply as hosting regular open calls for community leaders to join. Recent efforts such as FediForum, [^6] https://fediforum.org/ a virtual convening of Fediverse users in the style of “unconference” sessions, are now already undertaking such initiatives, reflecting the evolving nature of the Fediverse as a digital space.
At the same time, participants believed that technical infrastructuring could enable the kind of interactions and relationships between communities that must be in place for routines of inter-community governance to emerge. We walk through the imagined ecosystem to highlight three interdependent principles that would enable future design work to meet the challenges of coordinating governance across communities: modularity, polycentricity, and forkability.
@@ 6.1.1. Modularity: The Capacity for Parsimonious Communication
A community creates their Governance Nutrition Facts label through an interface through which it can cherry-pick modular choices for each dimension of community governance. This presumes that communities have a shared language about important governance decisions and how they are made, even if communities may vary in their selection. A modular system thus reduces the potential costs of communicating about community governance. Displaying a consistent set of dimensions and potential choices, modules allow for more parsimonious reading and review of community governance decisions by humans compared to raw textual description. Modules can also make community governance machine-readable. This can enable a new suite of automated tools around governance, e.g., identifying communities with similar practices, or alerting communities of potential governance frictions.
@@ 6.1.2. Polycentricity: The Capacity for Overlapping, Nuanced Relationships
  • Workshops noted that communities would benefit from getting information from one another and possibly sharing their moderation work. However, communities don’t trust all other communities equally, nor want to exchange information with everyone. A lack of infrastructure to coordinate with trusted others undermines their potential to share the burden of community governance.
@@
Communities create or join trust bubbles: sets of communities that they indicate a level of trust in, even if they are not all federated. In Figure 3, trust bubbles are represented by shaded blobs encompassing communities. Participants imagined that trust bubbles would help communities gather information about governance practices and issues efficiently and strategically. For example, communities could automatically share logs of their governance activity that is aggregated and shared in the “Issue Bulletin” for the trust bubble. Crucially, the bulletin is not a global repository; each trust bubble can instantiate its own Bulletin, and communities might navigate multiple Bulletins according to the various trust bubbles they are in. How each Bulletin processes and displays information differs based on the trust bubble’s trust levels (the “ladder of engagement”), which can vary over time. As with the modular system for community governance labels, trust bubbles have a modular set of choices for how to share information as well as engage in discussion. Trust bubbles might set processes to edit these choices as needed using existing tools such as PolicyKit by Zhang et al. (2020).
@@
Trust bubbles envision a polycentric [^8] For overview of polycentricity as a concept in the context of social media, see Jhaver et al. (2023) system of evolving coalitions that a community is a part of: multiple, overlapping sets of communities that coordinate with one another in different ways, for different reasons. Polycentricity is crucial because it enables a community to be in multiple trust bubbles that may or may not overlap; it also suggests multiple kinds of coordination across those trust bubbles. For example, a community may be in a smaller trust bubble that it actively shares raw moderation reports with, and a broader one that it looks to in order to track emerging issues (e.g., spam, scams, raids) on the network. In short, polycentricity involves supporting nuanced and diverse governing relationships between communities that can evolve over time.
@@
This dynamic and relational aspect of polycentricity in trust bubbles distinguishes it from existing ideas of how communities can ‘share the burden.’ For example, in the protocol for Bluesky (AT Protocol), “composable moderation” enables a community to label content (i.e., for filtering) independent of the app they use (Bluesky, 2023; Fraser and Kramer, 2025). Others can then subscribe to these labels, allowing them to avoid reinventing the wheel in terms of moderation. Trust bubbles imagine infrastructures that enable multiple communities to actively collaborate on such resources; different bubbles may reinvent the wheel as needed. Although we describe the Issue Bulletin to illustrate how this might manifest more concretely, our findings call more broadly for expanding how and when communities are able to support one another on governance problems.
@@ 6.1.3. Forkability: The Capacity for Variation
When a community generates their Governance Nutrition Facts label, they can fork a version from other communities and then adapt the decision(s) for each dimension of community governance. The options available in the forking system can draw from the Nutrition Facts of all communities in the network to provide a modular catalog of decisions to fork from along each dimension. It can also be filtered by trust bubble. In turn, communities can also contribute to this catalog, so that other communities might reference it in the future. The shared language that the modular forking system is based on is descriptive, not prescriptive. Along each dimension of community governance, the system prompts the community to reflect on whether the forked decision reflects their community’s actual values and goals.
@@
Forkability reduces the labor it takes to craft a set of governance practices and processes by allowing communities to avoid doing so from scratch. In practice, communities already copy and customize governance practices based on what they see in other communities (Tosch et al., 2024; Kiene et al., 2025), but are prone to path dependencies. We suggest that forkability also requires that communities have a fairly rich catalog of potential choices to copy from, as envisioned in the modular forking system in our ecosystem. However, forking is not just about copying; it is about adding, customizing, and discarding modular components as needed. Fundamentally about adaptation, forkability emphasizes making sure a community can tailor their governance practices to their independent needs. Communities can add their forks as new modules. Forkability thus also emphasizes that modular choices can evolve over time.
6.2. Implementing Infrastructures for Inter-Community Governance
We note that forkability alone does not ensure community autonomy. For example, communities may still face isomorphic pressures (DiMaggio and Powell, 1983; Caplan and boyd, 2018) to conform to “good” governance practices based on what they see in the Issue Bulletin, or given the trust bubbles they are a part of. However, polycentricity may help offset these pressures: the ecosystem described envisions that communities move across multiple trust bubbles and Issue Bulletins. Importantly, communities are able to coordinate and rely on these communities in nuanced ways. Rather than simply being ‘in or out,’ communities can maintain relationships with each of the bubbles in ways that help them maintain their own boundaries as a community.
@@ 6.2. Implementing Infrastructures for Inter-Community Governance
Indeed, modularity, polycentricity, and forkability are interdependent; the lack of one would undermine the others. Without modularity, communities lack the capacity for parsimonious communication that enables them to navigate nuanced relationships and distinguish themselves in forking. Without polycentricity, communities lack the capacity for diverse collaborative and cooperative relationships that allow them to assess modular options and know when to fork. Without forkability, communities lack the capacity to maintain their autonomy that sustain a diverse, polycentric system and to change the modular choices to fit their needs. All three principles must be attended to when implementing social and technical infrastructures for inter-community governance in design.
Our findings also suggest practical recommendations for the technical form implementations might take. Participants expressed hesitation about adopting new tools for the express purpose of inter-community governance, as well as noted how existing initiatives had not been as widely taken up as expected. Our findings suggest two reasons why. First, community leaders already felt like they were struggling to juggle a patchwork of tools for governing their communities. Second, the need for inter-community governance is not always salient. Rather, it emerges when a harm has already occurred due to a governance friction (e.g., norm violation). Both of these make on-boarding and logging onto a new system largely unappealing, producing yet another dimension of governance labor for communities.
We propose that infrastructures for inter-community governance be implemented via an independent but compatible layer of components that communities can integrate into their existing tools. This might involve creating private feeds that communities subscribe and contribute to, by defining a standard process for leveraging the existing protocol to create a private ‘meta’ server that trust bubble members are all federated to. The bulletins or even the modular forking system could simply be built as platform extensions, rather than standalone applications. One participant raised the idea of developing a “governance API’ that would define how all communities, regardless of software or platform, engage in inter-community governance. Seen another way, implementation of inter-community governance suggests devising an open protocol for coordinating governance questions, one that is orthogonal and complementary to the protocol that powers decentralized social media. This protocol may then catalyze a suite of tools that can be integrated into existing technologies in a community’s organizational routines.
6.2.1. Beyond Community-run Social Media
@@
The protocols that underpin systems such as the Fediverse tend to specify broad affordances for inter-community interactions, but provide little-to-no ways to support inter-community dynamics that would resolve tensions, reduce frictions, or help community leaders solve typical problems that arise. This has made the need for a layer of collective governance in this context particularly pressing. As interest in decentralizing systems through protocolization grows (Masnick, 2019; Kleppmann et al., 2024), we anticipate that other systems beyond the case of community-run social media will face similar questions and need to coordinate inter-community governance accordingly.
A remaining question is how ideas of inter-community governance translate beyond decentralized systems. Within centralized platforms, such mechanisms are likely less necessary because they are internalized by the entity operating the platform and whatever rules, interfaces, or dispute resolution mechanisms it sees fit to offer. However, we contend that the importance of community-level decision-making in those contexts (Seering, 2020; Jhaver et al., 2021b; Li et al., 2022c) keeps questions of inter-community governance pertinent. Communities in those contexts already do govern in relation to one another, shaped by the affordances of platform design, and would benefit from the kind of information-sharing and coalition-building imagined by inter-community governance (Kiene et al., 2025). On one hand, inter-community governance may inform changes to platform design, such as creating more back-end capacity for community organizers to coordinate or engage with one another. On the other hand, our findings also point to how communities might organize independently of a specific software. At a time when relationships between communities and platforms are in tension (Li et al., 2022a; Vertesi and Matias, 2023), building community capacities can offset the power of centralized platforms.
@@ 7. Limitations
Our findings are shaped by the subjectivity and self-presentation of our participants, as well as the interpretative nature of our qualitative work. We aimed to build a participant pool with a broad range of experiences on the Fediverse, but focused on individuals who were interested in governance questions for the purposes of our workshop. This allowed us to learn about perspectives of community organizers and developers, but future work evaluating the themes and ideas presented in this work with community members are likely to give insight into notions of fairness, transparency, and participation in the kinds of governance dynamics described in our findings. We also note that the participants involved in our workshops distinctly cared about governance. They had not only thought about practices and processes extensively, but also were willing to invest extra time, labor, and care in nurturing online spaces in ways that many others may not — even if they are community moderators. Insofar that decisions made in inter-community governance can pose an additional dimension of labor, future work must consider how to reduce barriers and costs to this kind of coordination work. In §6.2, we noted that inter-community governance is more useful when it is implemented as an infrastructural layer that provides access to resources and facilitates interaction across communities, rather than when it is implemented as standalone apps that require new sign-ups or technical learning. User studies with prototypes operationalizing some of the ideas described by participants would help offer more specific recommendations for designing such a layer, as well as identifying barriers to use. Finally, testing such prototypes outside the Fediverse will be important to assess the practical value and use of inter-community governance in a diverse array of contexts.

### v2

@@ 6.1. Modularity, Polycentricity, and Forkability: Principles for Inter-community Governance
Our findings suggest a collection of resources, norms, and spaces that can enable communities to collaborate and cooperate on governance issues from the bottom-up. In Figure 3, we sketch them as one ecosystem of inter-community governance on community-run social media. We present this illustration as a conceptual artifact that can prompt future design work aiming to support, shape, and scaffold dynamics of inter-community governance.
Some of the concepts presented above could be easily developed in the near future. However, as our findings note, the success of any system suggested in Figure 3 is likely to require shifts in broader social expectations around how communities operate on the Fediverse. For example, many workshop participants envisioned decentralized social media as consisting of “many fediverses,” an idea that they stated was considered controversial by those who want the benefits of scale. The specific tools we describe are not technical solutions, but aim to support norms of communities engaging with one another on governance questions when needed. We emphasize devising infrastructures that can create opportunities for communities to organize and reshape norms together. For example, a “nutrition label” alone could help a community communicate its practices to existing and potential members. However, a consistent, structured labeling schema for it would allow communities to engage in inter-community governance much more easily. To illustrate: as a machine-readable system, a repository of nutrition labels from a wide swathe of communities could enable people to quickly flag and review potentially incompatible communities (or find compatible ones).
Our perspective builds on prior work from HCI that intervenes in technical infrastructures that standardize processes (e.g., decision-making) or produce new possibilities (or constraints) (Edwards et al., 2010), as well as examines human infrastructures that do work to coordinate and manage these systems (Lee et al., 2006). Designing infrastructures presents opportunities for discussing social norms and discussing what types of decisions should be made at a global versus local level (Edwards et al., 2009). As we note in §5.6, this may start as simply as hosting regular open calls for community leaders. Recent efforts such as FediForum, [^6] https://fediforum.org/ a virtual convening of Fediverse users in the style of “unconference” sessions, are already undertaking such initiatives, reflecting the evolving nature of the Fediverse.
Technical infrastructuring can enable the kind of interactions and relationships between communities that must be in place for routines of inter-community governance to emerge. We walk through the imagined ecosystem to highlight three interdependent principles that would enable future design work to meet the challenges of coordinating governance across communities: modularity, polycentricity, and forkability.
@@ 6.1.1. Modularity: The Capacity for Parsimonious Communication
A community creates their Governance Nutrition Facts label via an interface through which it can cherry-pick modular choices for each dimension of community governance. This presumes that communities have a shared language about important governance decisions and how they are made, even if communities may vary in their selection. A modular system thus reduces the potential costs of communicating about community governance. Displaying a consistent set of dimensions and potential choices, modules allow for more parsimonious reading and review of community governance decisions by humans compared to raw textual description. Modules can also make community governance machine-readable. This can enable a new suite of automated tools around governance, e.g., identifying communities with similar practices, or alerting communities of potential governance frictions.
@@ 6.1.2. Polycentricity: The Capacity for Overlapping, Nuanced Relationships
  • Workshops noted that communities would benefit by getting information from one another and possibly sharing their moderation work. However, communities neither trust nor want to exchange information with each other equally. A lack of infrastructure to coordinate with trusted others undermines the potential to share the burden of community governance.
@@
Communities create or join trust bubbles: sets of communities that they indicate a level of trust in, even if they are not all federated. In Figure 3, trust bubbles are represented by shaded blobs encompassing communities. Participants imagined that trust bubbles would help communities gather information about governance practices and issues. For example, communities could automatically share logs of their governance activity that are aggregated in the “Issue Bulletin” for the trust bubble. Crucially, the bulletin is not a global repository; each trust bubble can instantiate its own Bulletin, and communities might navigate multiple Bulletins according to the various trust bubbles they are in. How each Bulletin processes and displays information differs based on the trust bubble’s trust levels, which can vary over time. As with the modular system for community governance labels, trust bubbles have a modular set of choices for how to share information as well as engage in discussion. Trust bubbles might set processes to edit these choices using existing tools such as PolicyKit by Zhang et al. (2020).
@@
Trust bubbles envision a polycentric [^8] For overview of polycentricity as a concept in the context of social media, see Jhaver et al. (2023a). system of evolving coalitions that a community is a part of: multiple, overlapping sets of communities that coordinate with one another in different ways, for different reasons. Polycentricity enables a community to be in multiple trust bubbles that may or may not overlap; it also suggests multiple kinds of coordination across those trust bubbles. For example, a community may be in a smaller trust bubble that actively shares raw moderation reports, and a broader one for tracking emerging issues (e.g., spam, scams, raids) on the network. In short, polycentricity involves supporting nuanced and diverse governing relationships between communities that evolve over time.
@@
This dynamic and relational aspect of polycentricity in trust bubbles distinguishes it from existing inter-community governance. For example, the AT protocol that underpins Bluesky affords “composable moderation” by enabling a community to label content (i.e., for filtering) independent of the app they use (Bluesky, 2023; Fraser and Kramer, 2025). Others can subscribe to these labels, allowing them to reduce their moderation burden. Trust bubbles imagine infrastructures that enable multiple communities to collaborate on such resources; different bubbles may design different solutions. The idea of the Issue Bulletin illustrates how this might manifest more concretely, as well as the potential benefits of expanding how and when communities are able to support one another on governance problems.
@@ 6.1.3. Forkability: The Capacity for Variation
When a community generates their Governance Nutrition Facts label, they can fork a version from other communities and adapt each dimension of community governance. The options available in the forking system can draw from the Nutrition Facts of all communities in the network to provide a modular catalog of decisions to fork from along each dimension. It can also be filtered by trust bubble. In turn, communities can also contribute to this catalog, so that other communities might reference it in the future. The shared language that the modular forking system is based on is descriptive, not prescriptive. Along each dimension of community governance, the system prompts the community to reflect on whether the forked decision reflects their values and goals.
@@
Forkability reduces the labor it takes to craft governance processes. Communities already copy and customize governance practices based on what they see in other communities (Tosch et al., 2024; Kiene et al., 2025), but are prone to path dependencies. We suggest that forkability also requires a heterogeneous catalog of potential choices to copy from, as envisioned in the modular forking system in our ecosystem. Forking is not just about copying; it is about adding, customizing, and discarding modular components as needed. Forkability ensures a community can tailor their governance practices to their independent needs. Communities can add their forks as new modules. Forkability thus also emphasizes that modular choices can evolve over time.
6.2. Tensions in Implementing Inter-Community Governance
@@
Modularity, polycentricity, and forkability are interdependent; the lack of one would undermine the others. Without modularity, communities lack the capacity for parsimonious communication that enables them to navigate nuanced relationships and distinguish themselves in forking. Without polycentricity, communities lack the capacity for diverse cooperative relationships that allow them to engage with one another, assess modular options, and know when to fork. Without forkability, communities lack the capacity to maintain their autonomy, sustain diversity, and evolve modular governance elements that fit their needs within a polycentric system. All three principles must be attended to when implementing infrastructures for inter-community governance in design.
The principles, and their interdependence, also reveal tensions and tradeoffs in inter-community governance. The emphasis on parsimonious communication in modularity underscores that the opacity and inconsistency in community governance makes it labor intensive to assess the good faith and legitimacy of communities online. The emphasis on evolving relationships with multiple bubbles of communities in polycentricity responds to concerns about potential re-centralization if communities begin to coalesce and coordinate. Likewise, the emphasis on autonomy and variation in forkability aims to avoid de facto re-centralization through institutional homogeneity.
Below we reflect on the two key tensions around labor and recentralization, and their implications for future work.
6.2.1. Labor
Even in a speculative mode, participants expressed hesitation about adopting inter-community governance tools. Community leaders already struggled to juggle a patchwork of tools for governing their communities. Inter-community governance was not always a priority. It became salient when a harm occurred due to a governance friction (e.g., norm violation). Under the circumstances, on-boarding, logging onto, and participating in a new governance system could seem unappealing, just another dimension of governance labor for communities. For example, crafting a tailored community nutrition label requires care and time, even if there are useful default modules that make communication more parsimonious. Likewise, an Issue Bulletin could produce a new source of information overload. Polycentricity suggests that communities could have many Bulletins, compounding these problems.
Attributes of our participants—all deeply invested in the work of community-building—may have led them to prefer tools that entail care and participation. Recall that many communities end up copying things like policies (Tosch et al., 2024; Kiene, 2024), and face risks of burnout (Schöpke-Gonzalez et al., 2024). Our participants were aware of broader trends and explicitly raised questions of feasibility, adoption, and sustainability. Future research should attend to the barriers to participating in these infrastructures as they are implemented. In the context of corporate-run platforms, failing to do so can risk enabling platform operators to exploit the care and labor of communities (Li et al., 2022a; Jhaver et al., 2023b; Matias, 2019b).
To reduce potential barriers, we propose that infrastructures be implemented via extensions rather than new standalone apps. The specific design of the tools described in Figure 3 will shape the labor associated with inter-community governance. Our findings suggest that inter-community governance could be leveraged to reduce the governance burdens communities face. This was most salient when discussing trust bubbles: participants imagined pooling moderation work with other communities. In a highly-trusted bubbles, communities might automatically share moderation decisions that are then applied by other communities in the bubble (e.g., deleting the same post, or flagging similar content), reducing the volume of things communities must independently deal with. Prior work in CSCW proposes similar ideas of leaning on assessments by trusted peers (Jahanbakhsh et al., 2022; Mahar et al., 2018). As Jahanbakhsh et al. (2022) find in the case of misinformation, preferences about the exact degree of automation in such systems (e.g., algorithmic filtering) vary widely. Future work testing prototypes should focus on these questions of labor, evaluating how to relieve the burdens of use — and toward strengthening the sustainability of communities.
6.2.2. Re-centralization
Communities may feel pressure to conform to “good” governance practices based on what they see in the Issue Bulletin, the trust bubbles they are a part of, or the community nutrition labels of large or influential communities. Participants raised concerns that trust bubbles could simply become spaces for re-centralization, or that nutrition labels could become prescriptive and lead to homogeneity. For example, making nutrition labels “forkable” (i.e., letting people copy and then customize from another community) could result in mindless copying or homogenization (Gaughan et al., 2025). Given the concerns around labor noted above, such risks seem real .
Communities often follow similar institutional patterns (Shaw and Hill, 2014) and organizational structures (Schneider, 2022, 2024), in addition to rules (Kiene, 2024). Some similarities may reflect shared goals or well-known best practices (Kraut and Resnick, 2012) — not all communities can be wildly different experiments in self-governance, and communities can learn from one another (Kiene et al., 2025). However, robust defaults can reduce barriers to entry, i.e., in creating a nutrition label that truly reflects a community’s goals. For example, even if a community forks a nutrition label, explicitly presenting a variety of choices about decision-making models or data retention policies can facilitate tailoring. We propose developing an open standard toolkit of modular “community governance choices” that covers a range of potential community configurations. Future UX and UI work can investigate how such toolkits balance between community autonomy and learning or copying.
Regardless, with an inter-community governance ecosystem, increased opportunities to interact about governance may increase isomorphic pressures (DiMaggio and Powell, 1983; Caplan and boyd, 2018) that push communities to make non-optimal decisions for themselves. Our participants suggested being able to move across multiple trust bubbles that hold distinct norms about what is shared, to whom. Thus, being ‘in’ a bubble can mean different things, enabling communities to have diverse types of relationships with other communities. Future work defining a taxonomy of trust to help communities navigate a “ladder of engagement” and underpin concepts like trust bubbles can shape how communities preserve autonomy amidst homogenizing pressures.
6.3. Beyond the Fediverse
@@
The protocols that underpin systems such as the Fediverse tend to specify broad affordances for inter-community interactions, but provide little-to-no ways to support inter-community dynamics, reduce frictions, or help community leaders solve typical problems. This makes the need for a layer of collective governance particularly pressing. However, interactions flow across communities in most social computing systems. We anticipate that other systems beyond federated, decentralized social media would benefit from inter-community governance.
Concepts like the community nutrition label or the trust bubbles evoke general questions about building and sustaining community. Thus, the ideas and design principles that underpin inter-community governance may be broadly applicable to a variety of social computing systems involving multiple communities, including lo-fi or indie ones. For example, a group starting a wiki for happenings in their city may want to connect with the leaders of other place-based communities to know how to deal with local emergencies, breaking news, and so on. Building the capacity to self-govern through ecosystems of inter-community governance can support communities to run their own systems, and we imagine that inter-community governance could support a diverse array of such systems.
Even in centralized platforms, community-level decision-making makes inter-community governance salient (Seering, 2020; Jhaver et al., 2021b; Li et al., 2022c). Communities in more centralized contexts already govern in relation to one another, and could benefit from the kind of information-sharing and coalition-building imagined by inter-community governance (Kiene et al., 2025). Inter-community governance may thus inform platform design, such as by justifying more back-end capacity for community organizers using Facebook Groups or Slack workspaces to coordinate with one another. Future work can also learn from the existing community governance features encoded in widely-used platforms. For example, Reddit provides a side panel with important information about the community for the platform: individual admins, community bookmarks, and community “flairs.” Although this does not necessarily reflect inter-community governance practices, empirical assessment of what is already made visible about community governance in popular systems can provide guidance on what still needs to be articulated.
@@ 7. Limitations
Our findings are shaped by the subjectivity and self-presentation of our participants, as well as the interpretative nature of our qualitative work. We aimed to build a participant pool with a broad range of experiences on the Fediverse, but focused on individuals who were interested in governance questions for the purposes of our workshop. This allowed us to learn about perspectives of community organizers and developers, but future work evaluating the themes and ideas presented in this work with community members are likely to give insight into notions of fairness, transparency, and participation in the kinds of governance dynamics described in our findings. Crucially, as we noted in 6.2, the participants in our workshops are heavily invested in governance. They had not only thought about practices and processes extensively, but also were willing to invest extra time, labor, and care in nurturing online spaces in ways that many others—even community moderators—may not. Future work should consider other perspectives and explore how to manage the potential barriers and costs to this kind of coordination work, particularly as we see growing calls to give more control to communities. In §6.2, we noted that inter-community governance is more useful when it is implemented as an infrastructural extension, rather than when it is implemented as standalone apps that require new sign-ups or technical learning. User studies with prototypes operationalizing some of the ideas described by participants would help offer more specific recommendations for designing such a layer, as well as identifying barriers to use. Finally, testing such prototypes outside the Fediverse will be important to refine our understanding of how and when to implement inter-community governance in a diverse array of contexts.